OpenStudy (anonymous):

How do you tell when an equation is an identity? How do you tell when an equation will have no solution? Give an example of both situations.

hero (hero):

An equation is an identity when it can be shown that both sides of an equation is equal for all x: For example 3(5x+2) = 15x + 4 15x + 4 = 15 x +4 If an equation has no solution, then there is no value that can be found for x and the expressions are not equal. For example: 2x + 5 = 2x - 7 Attempting to solve for x: 7 + 5 = 2x - 2x 12 = 0 False